# Broadening Discovery through Structural Models: Multimodal Combination of Local and Structural Properties for Predicting Chemical Features

### v1

@@ 5.2 QSAR tasks
MolCLR(GCN) [Wang et al., 2022] 85.4 0.0178 0.0317 3.25 1.41 0.95
MolCLR(GIN) [Wang et al., 2022] 91.6 0.0167 0.0225 2.88 1.25 0.65
ChemBERTa [Chithrananda et al., 2020] 177.2 - 0.0317 3.47 1.48 0.71
ECFP-BERT (ours) 159.4 0.0306 0.0148 2.09 1.03 0.81
BERT+GIN (ours) 83.4 0.0065 0.0060 0.35 0.27 0.31
BERT+GCN (ours) 84.5 0.0092 0.0078 0.55 0.36 0.54
BERT+Graphormer (ours) 81.9 0.0291 0.0142 2.32 1.001 0.90
Several of the utilized datasets—namely, BBBP [Sakiyama et al., 2021], Tox21 [Richard et al., 2020], ClinTox [Wu et al., 2018], BACE [Wu et al., 2018], MUV [Rohrer and Baumann, 2009], and HIV [Pan et al., 2007]—are specifically oriented towards classification tasks. We summarize the results for these datasets, along with comparisons to other proposed architectures, in Table 1. The receiver operating characteristic area under the curve (ROC-AUC) serves as the evaluation metric.
Conversely, the remaining datasets—QM7 [Blum and Reymond, 2009], [Rupp et al., 2012], QM8 [Ramakrishnan et al., 2015], QM9 [Ruddigkeit et al., 2012a], [Ruddigkeit et al., 2012b], FreeSolv [Mobley and Guthrie, 2014], ESOL [Delaney, 2004], and Lipo—focus on regression tasks. We utilize the mean absolute error (MAE) as the metric for the QM7, QM8, and QM9 datasets, while the mean squared error (MSE) serves as the metric for FreeSolv, ESOL, and Lipo. We present the findings and comparative analysis with other architectures in Table 2.
Most classification datasets are intrinsically linked to biochemical tasks and often feature relatively large molecules. Observations indicate that language models trained on SMILES representations, such as ChemBERTa, yield only modest performance metrics. This limitation arises from their inability to effectively account for atoms situated at significant distances from one another. In contrast, our models exhibit substantial improvements in metric performance when they incorporate a graph component, which enhances the capture of molecular structural information.
In the context of regression problems, where smaller molecules are more prevalent, language models demonstrate comparatively strong performance.
Thus, it becomes evident that both components of the architecture—namely, the graph and language models—are generally beneficial for achieving optimal model performance.
Graphormer, being a more complex model, tends to exhibit superior performance on large datasets. However, it often struggles with smaller datasets due to insufficient data for effective pre-training. Consequently, we recommend utilizing the BERT+GIN and BERT+GCN models for tasks characterized by limited data availability. Conversely, the BERT+Graphormer architecture is more suitable for intricate tasks that require establishing complex internal connections among nodes.
6 Conclusion
Our proposed set of architectures, consisting of ECFP RoBERTa (ECFP-BERT) and bimodal configurations that integrate ECFP-BERT as the language branch alongside Graph Convolutional Network (GCN), Graph Isomorphism Network (GIN), or Graphormer as the graph branch, has demonstrated some of the most promising performance metrics compared to existing models in the domain for various quantitative structure-activity relationship (QSAR) problems across a range of well-established benchmarks.
While specialized descriptors typically outperform transformer-based models for these challenges, these benchmarks serve as a simplified context, thereby allowing us to assess the quality and versatility of our architecture without the confounding influence of large-scale superstructures commonly encountered in more complex scenarios.
To further validate our model, we intend to explore additional challenges, including co-crystal prediction, the prediction of nuclear magnetic resonance (NMR) spectra from molecular structure, and other physicochemical tasks. For such demanding tasks, transformer-based architectures often yield significantly superior results compared to straightforward augmentations of task-specific descriptors.
However, it is important to note that addressing these tasks will necessitate considerable modifications to the existing architecture. Consequently, the outcomes will be contingent not only upon the quality of the embeddings currently utilized but also on the enhancements made to the model architecture itself.
Appendix A ECFP construction algorithm
ECFP [Rogers and Hahn, 2010] is a so-called circular fingerprint that assigns a two-dimensional hash array to each molecule using the following algorithm:
  1. 1.
    The initial step is assigning an integer identifier to each atom.
  2. 2.
    The iterative update stage, in which the identifier of each atom is updated with the identifiers of its neighbours.
  3. 3.
    Duplicate removal - a stage in which several occurrences of the same feature are reduced to a single representation in the feature list.
One iteration for a single atom is as follows:
  1. 1.
    An array of integers containing the iteration number and the ID of the given atom is initialized.
  2. 2.
    The attached atoms are sorted in deterministic order using the bond order (single, double, triple, and aromatic) and the current ID of each attached atom. or each attachment, the attachment ID and bond order are added to the array.
  3. 3.
    The array is hashed into a single 32-bit integer. This is the new atom identifier.
Appendix B Graph models
@@
Appendix C Some training details
@@
Appendix D Testing datasets (QSAR)
@@
The QM7 dataset is a curated subset of GDB-13, a comprehensive database containing nearly one billion stable and synthetically accessible organic molecules. Specifically, QM7 includes 7,165 molecules, each composed of up to 23 atoms, with a focus on seven heavy atoms: carbon (C), nitrogen (N), oxygen (O), and sulfur (S). This dataset not only provides a diverse array of molecular structures—such as double and triple bonds, cyclic compounds, carboxylic acids, cyanides, amides, alcohols, and epoxides—but also features the Coulomb matrix representation of these molecules. Additionally, the atomization energies for the QM7 molecules are computed using methods aligned with the FHI-AIMS implementation of the Perdew-Burke-Ernzerhof hybrid functional (PBE0).

### v2

@@ 5.2 QSAR tasks
MolCLR(GCN) [Wang et al., 2022] 85.4±2.7 0.0178±0.0003 0.0317±0.0005 3.259±0.261 1.419±0.040 0.957±0.010
MolCLR(GIN) [Wang et al., 2022] 91.6±3.1 0.0167±0.0004 0.0225±0.0003 2.884±0.249 1.253±0.037 0.651±0.004
ChemBERTa [Chithrananda et al., 2020] 177.2±4.0 - 0.0317±0.0005 3.471±0.085 1.487±0.107 0.721±0.003
Uni-Mol [Zhou et al., 2023] 41.8±0.2 0.0156±0.0001 0.0132±0.0003 1.480±0.048 0.788±0.024 0.603±0.010
GEM [Fang et al., 2022] 58.9±0.8 0.0171±0.0001 0.0246±0.0003 1.877±0.094 0.798±0.029 0.660±0.008
GROVER (base) [Rong et al., 2020] 94.5±3.8 0.0218±0.0004 0.0197±0.0003 2.186±0.052 0.983±0.090 0.817±0.008
GROVER (large) [Rong et al., 2020] 92.0±0.9 0.0224±0.0003 0.0186±0.0025 2.272±0.051 0.895±0.017 0.823±0.001
Molformer [Wu et al., 2023] 55.2±0.8 0.0095±0.0005 0.0139±0.0004 - - -
MolFromer-XL [Ross et al., 2022] - 0.0102±0.0002 0.0164±0.0002 0.571±0.027 0.290±0.011 0.551±0.002
SubD-BERT (ours) 55.4±1.3 0.0126±0.0003 0.0148±0.0003 0.859±0.069 0.292±0.013 0.514±0.003
BERT+GIN (ours) 49.8±1.1 0.0083±0.0002 0.0145±0.0003 0.530±0.041 0.331±0.018 0.526±0.009
BERT+GCN (ours) 50.1±1.7 0.0098±0.0003 0.0166±0.0004 0.731±0.098 0.357±0.026 0.540±0.016
BERT+Graphormer (ours) 40.6±0.9 0.0114±0.0003 0.0134±0.0003 0.823±0.091 0.291±0.023 0.589±0.010
XGBoost (descriptors, ours) 69.2 0.0208 0.0174 5.040 1.233 1.013
LightGBM (descriptors, ours) 74.1 0.0203 0.0173 5.435 1.217 0.997
SVR (descriptors, ours) 143.4 0.0301 0.0334 6.207 1.830 1.171
The classification benchmark datasets comprised BBBP [Sakiyama et al., 2021], Tox21 [Richard et al., 2020], ClinTox [Wu et al., 2018], BACE [Wu et al., 2018], MUV [Rohrer and Baumann, 2009], and HIV [Pan et al., 2007]. Performance metrics, summarized in Table 1, utilize the receiver operating characteristic area under the curve (ROC-AUC) (Table 2). The regression benchmarks included QM7 [Blum and Reymond, 2009, Rupp et al., 2012], QM8 [Ramakrishnan et al., 2015], QM9 [Ruddigkeit et al., 2012a, b], FreeSolv [Mobley and Guthrie, 2014], ESOL [Delaney, 2004], and Lipo. Evaluation metrics consisted of mean absolute error (MAE) for QM7, QM8, and QM9, and mean squared error (MSE) for FreeSolv, ESOL, and Lipo (Table 2).Three runs were conducted for each dataset, and the experiments were performed on 4 A100 GPUs.
Notably, the classification datasets predominantly involve biochemical tasks with relatively large molecules. Language models trained on SMILES representations, such as ChemBERTa, demonstrate limited efficacy in these cases, likely due to their inability to adequately capture long-range atomic interactions. In contrast, our approach, which decomposes molecules into substructures, exhibits marked improvements when supplemented with graph-based components that effectively encode structural information.
Conversely, for regression tasks (in datasets such as QM8 [Ramakrishnan et al., 2015] and QM9 [Ruddigkeit et al., 2012a]) focusing on smaller molecules, language models achieve comparatively poor results individually. The number of substructures in such molecules is small, and the compounds are represented by a limited number of substructures and descriptors, which implies insufficient data. Taken together, these findings underscore the complementary contributions of both graph and language model components in optimizing predictive performance.
Graphormer, as a complex model, generally performs better on large datasets but struggles with smaller ones due to limited training data. Therefore, we recommend BERT+GIN and BERT+GCN for tasks with limited data, while BERT+Graphormer is better suited for complex tasks that require intricate node relationships.
For the general case, our models outperform other frameworks, including MolFormer-XL [Ross et al., 2022], which was trained on a 1.1 billion size dataset for approximately 3200 compute hours on Nvidia V100 GPU.
6 Limitations and further improvements
Limitations. In chemoinformatics, we frequently encounter challenges that require building predictive models from very limited training datasets, typically consisting of only 100–200 samples. In these scenarios, our models tend to be overly complex and are frequently outperformed by simpler models trained on task-specific descriptors tailored for the problem at hand.
Furthermore, similar difficulties arise in cases where the training sample size is relatively small (on the order of a few thousand samples), but the compounds under study exhibit a substantial domain shift compared to our training set, usually a subset of the PubChem database. This issue is particularly pronounced when working with inorganic compounds or polymers. Traditional pre-training approaches offer limited benefit here, as BRICS decomposition is primarily designed for organic molecules structurally similar to drug-like compounds, and its fragmentation rules require adaptation to effectively handle these chemically distinct classes.
Future works. Further improvements involve training the models on a larger data sample: approximately 100 million molecules or more. Given that the current approach outperforms existing methods with a rather modest training sample, this should give a significant performance gain.
Appendix A Descriptors
A.1 General description
The interplay between various molecular properties has long been recognized, and exploiting these relationships to predict molecular behavior is a standard practice in chemoinformatics [Hansch et al., 1962, 1963, Hansch and Fujita, 1964]. Central to this approach is the use of molecular descriptors – quantitative representations of molecular structure and characteristics – that are typically grouped into six categories. Substructural descriptors, such as those in fingerprinting methods like MACCS keys or PubChem fingerprints, encode the presence or absence of specific structural motifs as bit vectors. Topological descriptors derive from the molecule’s two-dimensional graph, capturing connectivity patterns through graph invariants. Geometric descriptors describe the three-dimensional shape and spatial configuration by considering molecular conformations or interatomic distances. Electronic descriptors relate to electron distribution and chemical reactivity, using parameters like orbital energies and atomic charges. Physico-chemical descriptors quantify bulk properties, including polarity, solubility, and hydrophobicity. Hybrid descriptors – like circular fingerprints – combine aspects of substructure and local topology to generate fixed-length numerical representations tailored to computational modeling.
Among topological descriptors, the Wiener index is particularly prominent. It measures molecular branching by summing shortest path distances between all atom pairs, originally for hydrocarbons but now used across diverse chemical applications such as predicting boiling points and drug-likeness [Wiener, 1947]. Another key descriptor is the octanol-water partition coefficient (logP), which reflects hydrophobicity through a molecule’s equilibrium distribution between octanol and water phases. LogP is critical in drug design because it affects membrane permeability, solubility, and bioavailability [Estrada, 2008].
For 3D representations, the Universal Force Field (UFF) energy estimates steric strain and conformational stability by approximating atomic interactions using simplified force fields. Although less accurate than quantum mechanical calculations, UFF energy is computationally efficient for large-scale virtual screening [Rappé et al., 1992]. Additionally, ring descriptors – especially distinguishing aromatic from non-aromatic rings – are vital in characterizing molecular stability, electron delocalization, and reactivity. Aromatic rings, with their conjugated $\pi$-systems, often enhance binding affinity in drug-receptor interactions, whereas non-aromatic (aliphatic) rings contribute to structural rigidity and three-dimensional shape.
These descriptors form the foundation for quantitative structure – activity and structure–property relationship (QSAR/QSPR) models, which predict molecular properties from structural data. Traditional QSAR approaches typically compute descriptors for the whole molecule, which works well for simpler compounds but may overlook subtle influences of individual substructures, especially in larger or multifunctional molecules [Estrada, 2008]. To overcome this, recent strategies focus on fragment-level descriptors – generating features for discrete molecular fragments rather than the entire molecule. This fragment-based modeling enhances interpretability and allows more precise predictions by capturing localized structural effects, thereby improving our understanding of complex molecular behavior.
When employed as input representations for language models, molecular descriptors hold significant promise. Some descriptors possess positional features similar to natural language sequences while also encoding essential physicochemical properties of molecules. This dual nature allows language models to capture not only statistical patterns but also meaningful chemical relationships, thereby improving their capability to derive insights that are relevant to molecular behavior.
A.2 Specific for our work
For this study, we computed a set of molecular descriptors for each substructure to serve as input features for the language model (LM). These descriptors comprehensively capture aspects of atomic composition, bonding patterns, topological indices, and physicochemical properties, providing the LM with rich, chemically meaningful representations.
Atomic Composition
[…]
Substructure Size
  • Total atom count within the substructure (single value).
Bonding Patterns
  • Counts of various bond types: single, double, triple, and aromatic, represented as four discrete values.
  • Total number of bonds in the substructure (single value).
Topological Descriptors
  • Wiener index: Computed as the sum of shortest topological distances between all pairs of atoms within the substructure, serving as a quantitative measure of molecular branching Wiener [1947].
Physicochemical Properties
  • logP (octanol-water partition coefficient): Classified into seven discrete bins reflecting hydrophobicity levels, as detailed in Table 3. This categorization facilitates interpretability and relevance to biological contexts.
    Category logP Range Description
    1 $<-2.0$ Extreme hydrophilic
    2 $-2.0\leq\text{logP}<-0.5$ Strong hydrophilic
    3 $-0.5\leq\text{logP}<0.0$ Moderate hydrophilic
    4 $0$ logP < 1.0 Neutral
    5 $1.0\leq\text{logP}<2.0$ Moderate hydrophobic
    6 $2.0\leq\text{logP}<4.0$ Strong hydrophobic
    7 $\geq 4.0$ Extreme hydrophobic
    Table 3: logP categorization scheme
  • Universal Force Field (UFF) energy: Categorized into seven discrete ranges representing steric strain and conformational stability; see Table 4. Despite its approximate nature compared to quantum calculations, UFF energy offers a computationally efficient descriptor of molecular geometry.
    Category Energy Range (kcal/mol) Description
    1 $<0$ Stable complex
    2 $500\leq E<$ 50 Very stable
    3 $10050\leq E<$ 100 Stable
    4 $150100\leq E<$ 150 Moderate
    5 $200150\leq E<$ 200 Unstable
    6 $300200\leq E<$ 300 High energy
    7 $\geq 300$ Extreme energy
    Table 4: UFF energy categorization scheme
Ring Systems
  • Number of atoms that participate in at least one ring (single value).
  • Number of complete rings fully contained within the substructure (single value).
In summary, this results in a feature vector of dimension 23 for each molecular substructure, integrating diverse structural and physicochemical information suitable for language model input.
Appendix B BRICS Decomposition Methodology
The Breaking of Retrosynthetically Interesting Chemical Substructures (BRICS) method [Degen et al., 2008] offers a systematic, rule-based framework for fragmenting molecules into chemically meaningful components that correspond to synthetic building blocks. This fragmentation strategy is founded on the principles of retrosynthetic analysis [Corey, 1991], wherein complex molecules are iteratively broken down into simpler precursors by cleaving bonds commonly formed or targeted in synthetic reactions.
B.1 BRICS Cleavage Rules
The BRICS algorithm applies a set of 16 predefined bond cleavage rules designed to target specific chemical environments. These rules selectively break bonds adjacent to key atoms, ensuring the preservation of functional group integrity and chemical relevance. The full repertoire of bonds eligible for cleavage includes:
  • Single bonds between aliphatic carbon atoms and any of the following heteroatoms:
[…]
      C-N bonds (Rule 1)
[…]
      C-O bonds (Rule 2)
[…]
      C-S bonds (Rule 3)
[…]
      C-P bonds (Rule 4)
  • Bonds in cyclic systems:
[…]
      C=C in conjugated systems (Rule 5)
[…]
  • Bonds adjacent to carbonyl groups:
[…]
      C-C(=O) (amide bonds, Rule 9)
[…]
      N-C(=O) (peptide bonds, Rule 10)
[…]
      O-C(=O) (ester bonds, Rule 11)
  • Specialized cleavages:
[…]
      C#C triple bonds (Rule 12)
[…]
      C-Si bonds (Rule 13)
[…]
      S(=O)-N sulfonamide bonds (Rule 14)
[…]
      C-B boronic ester bonds (Rule 15)
[…]
      C-Sn stannane bonds (Rule 16)
Each cleavage generates molecular fragments that incorporate dummy atoms at the original attachment sites, thereby retaining critical information about potential points for recombination. These cleavage rules are applied recursively until no additional bonds meeting the criteria remain, resulting in a comprehensive set of BRICS fragments.
B.2 Chemical Rationale
The choice of targeted bond types is grounded in their widespread occurrence in synthetic organic chemistry [Carey and Sundberg, 2007] and pharmaceutical molecules [Kumar et al., 2012]. This targeted fragmentation approach offers several key advantages, including:
  • Preservation of essential pharmacophoric features within the generated fragments, ensuring retention of bioactive characteristics.
  • Production of building blocks that are synthetically accessible, facilitating practical chemical synthesis.
  • Maintenance of chemically valid valency states in all fragments, preserving structural integrity and realism.
  • Compatibility with combinatorial library design workflows, enabling efficient exploration of chemical space [Kodadek, 2011].
B.3 Application in Descriptor Calculation
In our preprocessing pipeline, molecules are first decomposed into substructures using the BRICS methodology, after which molecular descriptors are computed for each fragment individually. This fragment-based approach provides several distinct analytical advantages:
  • Enables more precise characterization of local chemical environments that predominantly influence specific molecular properties.
  • Facilitates enhanced interpretability by allowing property attributions at the fragment level.
  • Improves the treatment of structurally complex molecules, where different subregions may exert contrasting effects on target properties.
  • Aligns naturally with fragment-based drug design strategies, promoting integration with established pharmaceutical workflows [Bian and Xie, 2018].
The fragment-centric representation is consistent with the concept of molecular signatures, in which molecular properties arise from both additive and nonlinear interactions among constituent substructures. By individually evaluating these fragments, our model is capable of pinpointing key structural motifs that significantly impact the target properties, while simultaneously ensuring synthetic feasibility through the use of the BRICS framework.
Appendix C Graph models
@@
Appendix D Some training details
@@
Appendix E Testing datasets (QSAR)
@@
The QM7 dataset is a curated subset of GDB-13, a comprehensive database containing nearly one billion stable and synthetically accessible organic molecules. Specifically, QM7 includes 7,165 molecules, each composed of up to 23 atoms, with a focus on seven heavy atoms: carbon (C), nitrogen (N), oxygen (O), and sulfur (S). This dataset not only provides a diverse array of molecular structures – such as double and triple bonds, cyclic compounds, carboxylic acids, cyanides, amides, alcohols, and epoxides – but also features the Coulomb matrix representation of these molecules. Additionally, the atomization energies for the QM7 molecules are computed using methods aligned with the FHI-AIMS implementation of the Perdew-Burke-Ernzerhof hybrid functional (PBE0).
@@
Appendix F Code
Our code is available by link  [^1] Our code for all experiments is accessible on https://anonymous.4open.science/r/thinking-like-a-chemist-EC7B..